# Vortices and Factorization

### v2

@@ 7 KWCC Transformation and Terminating Configurations
It is worth mentioning that all families of equilibrium patterns obtained above are related to the following chain of transformations that starts from the free Schrodinger operator
$A_{-1}A_{-1}^{*}\xrightarrow{\gamma_{0}}A_{0}^{*}A_{0}\to A_{0}A_{0}^{*}\xrightarrow{\gamma_{1}}A_{1}^{*}A_{1}\to\dots,\quad\gamma_{i}\in\{1/2,1,2\},$
where $A_{-1}=-A^{*}_{-1}=\partial_{z}$, and $\xrightarrow{\gamma_{i}}$ stand for exponentiations of kernels. For example, for the even bi-spectral family $\gamma_{0}=1/2$ and $\gamma_{i}=1$, $i>0$. Chains of this type are generalizations of (35).
@@ 8 Translating Configurations and Baker-Akhieser Functions
Translating configuration of vortices are solutions of (8). To generate them we consider the following eigenfunction
$\psi(k,z)=C\prod_{i=1}^{N}(z-z_{i}(k))^{{\cal Q}_{i}}e^{kz},$ (67)
of the Schrodinger operator
@@
The condition that potential $u=\psi^{\prime\prime}/\psi-k^{2}$ does not have simple poles (i.e. it is of form (48)) is equivalent to the equilibrium conditions (8) for a translating configuration. This can be shown by arguments similar to that used for the static configurations in Section 5.
In the case of the factorization chain (35) starting with the free Schrodinger operator, there exists an eigenfunction, called the “rational Baker-Akhieser function” [^10] In theory of integrable hierarchies of non-linear PDEs, the Baker-Akhieser function is a solution of an infinite system of linear differential equations. Associated integrable hierarchy arises as a condition of compatibility of the system. Here we deal with the Baker-Akhieser function of the KdV hierarchy., which has the form (67). It is an eigenfunction of $H_{j}$ that is obtained by the action of an intertwining operator with rational coefficients on the eigenfunction $e^{kz}$ of the free Schrodinger operator, i.e. $\Psi=C(k)T_{j}[e^{kz}]$. Since the $z$-independent factor $C(k)$ does not play role in our considerations, we will use the eigenfunction
$\psi_{j}=T_{j}[e^{kz}]=\frac{q_{j}}{p_{j}}e^{kz},$ (68)
where $p_{j}$ and $q_{j}$ are polynomials in $z$. As we have shown in section 6, the corresponding potential $u_{j}$, obtained by $j$ Darboux transformations from $u_{0}=0$, is of the form (48). Since $u_{j}$ does not contain simple poles, eigenfunction (68) corresponds to a translating configuration. It will be shown in section 12 that the class of such configurations contains all possible translating patterns consisting of finite number of vortices with ${\cal Q}_{i}=\pm 1$.
A constructive way to write down $\psi_{j}$ is to use the Wronskian representation of $T_{j}$, given by (44, 45). Then, from (68) we get
@@
$q_{j}=e^{-kz}{\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j},e^{kz}\right],\quad p_{j}={\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j}\right]=P_{j},$ (69)
where $P_{j}$ is the $j$th Adler-Moser polynomial. When $k\not=0$, degree of $q_{j}(z)$ equals to that of $p_{j}(z)$. In the electrostatic interpretation (see section 1), this means that the total charge has to be zero $\sum_{i}{\cal Q}_{i}=0$ in order the system to be at rest in the homogeneous electric field $k\not=0$. In the $k\to 0$ limit
@@ 9 Equilibria in Non-Homogenous Background Fields
$\psi(z)=C\prod_{i=1}^{N}(z-z_{i})^{{\cal Q}_{i}}e^{\Phi(z)},\quad H\psi=\lambda\psi,$ (70)
@@
Equations (71) are conditions of equilibrium of charges in an external harmonic electric field [^11] Such conditions were first presented by Stieltjes [36, 37] for roots of orthogonal polynomials. or equilibrium of vortices in a background flow, see section 1. Function $\Phi(z)$ is a potential of the external field/background flow. Provided (71) holds,
@@
$\tilde{\psi}(z)=C\prod_{i=1}^{\tilde{N}}(z-\tilde{z}_{i})^{{\cal\tilde{Q}}_{i}}e^{\Phi(z)},\quad H\tilde{\psi}=\eta\tilde{\psi},$
@@
will be also quasi-rational of the form similar to (70), i.e.
@@
does not have first-order poles, $\hat{\psi}$ corresponds to a new equilibrium configuration in the same external field. For a set of eigenfunctions of $H$ (“seed functions”) of the form
$\varkappa_{i}=R_{i}(z)e^{\Phi(z)},\quad H\varkappa_{i}=\lambda_{i}\psi_{i},$
where $R_{i}(z)$ are rational in $z$, we can apply several Darboux transformations resulting in various equilibrium configurations. According to the Darboux-Crum theorem (38,39), the results of compositions of $n$ transformations can be written in the Wronskian form
$\psi^{(n)}(z)=\frac{{\cal W}[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n-1}},\varkappa_{k_{n}}]}{{\cal W}[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n-1}}]}.$ (73)
For example, for harmonic oscillator $u=z^{2}/2$, $\Phi(z)=-z^{2}/2$ and $R_{n}$ are Hermite polynomials. In this case, transforms (73) correspond to various equilibria of charges in the external quadrupole field or vortex patterns in vicinity of the stagnation point of the flow encountering a solid boundary (for details and other examples, see e.g. [14, 25]). Note, that in difference from the systems considered in the previous sections, configurations corresponding to Hermite polynomials do not depend non-trivially on continuous parameters and are rather parametrised by sequences of distinct integers $k_{1}<k_{2}<\dots<k_{n}$.
@@ 10 Equilibria of Vortex Streets - Odd Family
A periodic vortex street is a vortex pattern consisting of an infinite number of vortices situated periodically along some direction. A vortex street can be equivalently considered as a configuration of $N$ vortices (or charges) on a cylinder. In what follows by the terms “vortex street”, “periodic configuration” or “periodic pattern” we mean configuration of a finite number of charges on cylinder. Without loss of generality we can set the translation period to $\pi$. Since $\sum_{n=-\infty}^{\infty}\frac{1}{z+\pi n}=\cot z$, the stationarity condition (5) for the translating (with a speed ${\rm i}\bar{k}/(2\pi)$) vortex street becomes
@@
and for static periodic patterns (see (6)) we have
@@
Similarly to vortex patterns on the plane that are related to the rational solutions of the KdV hierarchy or the odd bi-spectral family, the trigonometric soliton solutions of the hierarchy are related to the patterns on a cylinder [19, 21, 25]. The latter can be shown by methods similar to those used in sections 5, 8.
@@
where ${\cal Q}_{i}$, $z_{i}$ correspond to the periodic static configuration (75). The “trigonometric-rational” here means that all ${\cal Q}_{i}$ are integers. Trigonometric-rational are special cases of the trigonometric-factorizable functions, where ${\cal Q}_{i}$ are real. Function (76) is an eigenfunction of the Schrodinger operator
@@
similarly to the rational case, we come to the conclusion that the static equilibrium condition is equivalent to absence of first-order poles in the potential and
@@
A family of the trigonometric-rational configurations can be generated by the chain of Darboux transformations (34) with $\eta_{i}=k_{i}^{2}$, where $0<k_{1}<k_{2}<\dots<k_{i}<\dots$ is a sequence of positive integers. The chain starts from the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$ and it is constructed with the help of “seed functions” that are its periodic eigenfunctions
@@
Then, according to the Crum theorem (38,39), at $n$th step of the chain we have
@@
Obviously, $\tau_{n}(z)$ are trigonometric polynomials, $\psi^{(n)}(z)$ are trigonometric-rational functions and corresponding potentials
@@
are of the form (77) and do not have simple poles. Thus, eigenfunctions (79) correspond to the static vortex street configurations. Note that potentials (80) are periodic $n$-soliton solutions of the KdV hierarchy and $\tau_{n}$ are its $n$-soliton $\tau$-functions (see e.g. [29] and references therein). We will demonstrate in section 12 that the set of these configurations actually contains all static vortex street patterns with ${\cal Q}_{i}=\pm 1$.
For translating configurations (see (74)) related to the soliton solutions we can apply metod similar to that used in section 8. By analogy with quasi-rational (67), we now consider the trigonometric quasi-rational eigenfunctions
@@
According to the Darboux-Crum theorem (38), eigenfunctions, obtained by a sequence of the Darboux transformation from the formal eigenfunction $e^{kz}$ of the free Schrodinger operator, equal
@@
with $\varkappa_{k_{i}}$ given by (78). Obviously $\psi^{(n)}(k,z)$ are of the form (81) and therefore correspond to the translating vortex street configurations. Since each $\varkappa_{k_{i}}$ (78) depends on a free parameter, the translating and static configurations related to soliton solutions depend on $n$ free parameters as well as on the sequence of integers $0<k_{1}<k_{2}\dots<k_{n}$. Up to a $z$-independent common factor $C(k)$, (82) is the trigonometric Baker-Akhieser function of the KdV hierarchy [^12] It is also the “angular part” of the Baker-Akhieser function of 2d quantum algebraically integrable systems, see e.g. [5],[12] and references therein..
The class of configurations determined by (82,78) contains all possible translating vortex street patterns with ${\cal Q}_{i}=\pm 1$, as will be shown in section 12.

### v3

@@ 7 KWCC Transformation and Terminating Configurations
8 $\Lambda=2$ Case and Third-Order Operators
In the $\Lambda=1$ case, the KWCC transformation (64) is a Darboux transformation at the zero energy level (57), since here $\gamma_{i}=1$. However, when $\Lambda=2$, i.e., $\gamma_{i}\in\{1/2,2\}$, the KWCC transformation is a composition of the Darboux transformation and exponentiation. This raises a natural question about the existence of a Darboux transformation in the $\Lambda=2$ case. The answer to this question is affirmative. However, unlike the $\Lambda=1$ case, one must consider Darboux transformations of third-order differential operators rather than second-order Schrödinger operators [35].
In more detail: The $\Lambda=2$ specification of the bilinear equation (10) can be written in the Schrodinger form as
$\left(-\partial_{z}^{2}-6(\log q)^{\prime\prime}\right)\left[\frac{p}{q^{2}}\right]=0.$
As follows from Section 3, for $q=q_{n}$, two linearly independent solutions of the above equation are $p_{n-1}/q_{n}^{2}$ and $p_{n}/q_{n}^{2}$ respectively. Since a linear combination of $p_{n}$ and $p_{n-1}$ is just a reparametrized $p_{n}$ or $p_{n-1}$, we can write
$\left(-\partial_{z}^{2}+u_{n}\right)\phi=0,\quad u_{n}=-6(\log q_{n})^{\prime\prime},$ (72)
where
$\phi=p_{n}/q_{n}^{2}\quad{\rm or}\quad\phi=p_{n-1}/q_{n}^{2}.$ (73)
In contrast to the $\Lambda=1$ case, the above sequence of second-order operators cannot be generated by Darboux transformations (31), because now $u_{n\pm 1}\not=u_{n}-2(\log\phi)^{\prime\prime}$. To proceed, we note that according to the first-order recurrence relations (25), $p_{n}$ in (72, 73) can be replaced with $q_{n+1}^{\prime}q_{n}-q_{n+1}q_{n}^{\prime}$, so $\phi=(q_{n\pm 1}/q_{n})^{\prime}$. Then, from (72) it follows that the ratio $q_{n\pm 1}/q_{n}$ satisfies the third-order differential equation
$L\left[\frac{q_{n\pm 1}}{q_{n}}\right]=0,\quad L=\partial_{z}^{3}-u_{n}\partial_{z}.$
The general solution $\varkappa$ of this differential equation, i.e., the general solution of
$L\varkappa=0,\quad L=\partial_{z}^{3}-u_{n}\partial_{z},$ (74)
is a linear combination of three functions: $\left\{q_{n+1}/q_{n},q_{n-1}/q_{n},1\right\}$ These functions are linearly independent because $q_{n-1},q_{n},q_{n+1}$ have distinct degrees.
Any non-constant solution of (74), i.e., $\varkappa=q/q_{n}$, where $q=C_{1}q_{n+1}+C_{0}q_{n}+C_{-1}q_{n-1}$, corresponds to an equilibrium configuration, since $\phi=\varkappa^{\prime}=p/q_{n}$, with $p=q^{\prime}q_{n}-q_{n}^{\prime}q$ being a polynomial. The polynomial $q=C_{1}q_{n+1}+C_{0}q_{n}+C_{-1}q_{n-1}$ is a reparametrized $q_{n\pm 1}$ or reparametrized $q_{n}$.
Thus, the set $\{C_{n}^{\pm}q_{n\pm 1}/q_{n}$: $n\in\mathbb{Z}\}$, where $C_{n}^{\pm}$ are arbitrary constants, contains all solutions [^18] The constant solution corresponds to the limit $C_{n}^{\pm}\to 0$, $s_{\pm n}\to\infty$, $C_{\pm n}^{\pm}s_{\pm n}={\rm const}$, $n>0$ (see (28)). of (74). It corresponds to a complete set of equilibrium configurations of two species of charges with ${\cal Q}_{i}\in\{-1,2\}$. The zeros of $\phi=\varkappa^{\prime}=(q_{n\pm 1}/q_{n})^{\prime}$ correspond to the positions of charges with ${\cal Q}_{i}=-1$, while its poles correspond to positions of charges with ${\cal Q}_{i}=2$.
9 Darboux Transformation for Third-Order Operators and Equilibrium Configurations
Darboux transformations for third-order operators of the form $\partial_{z}^{3}-u\partial_{z}$ were found by Aiyer et al. [8]. Their derivation, via factorization, was presented by Athorne and Nimmo in [7]. Here, we will derive a zero-level chain of Darboux transformations for such operators.
We consider the third-order operator $L$ and its transform $\hat{L}$
$L=\partial_{z}^{3}-u\partial_{z},\quad\hat{L}=\partial_{z}^{3}-\hat{u}\partial_{z}.$ (75)
Let $\varkappa$ be a non-constant element of kernel of $L$. Operator $L$ can be expressed as the product of the second and the first-order factors
$L=BA,$ (76)
where
$B=\partial_{z}^{2}+v\partial_{z}-v^{\prime}-\frac{v^{\prime\prime}}{v},\quad A=\partial_{z}-v,\quad v=\frac{\varkappa^{\prime}}{\varkappa}.$ (77)
The potential $u$ is expressed in terms of $v$ as follows
$u=3v^{\prime}+v^{2}+\frac{v^{\prime\prime}}{v}.$ (78)
In contrast to the case of the Schrodinger operator, permutation of factors in (76) maps $L$ into $\hat{L}$ of the similar type (i.e. operator containing only third and first derivative, as in (75)), only if [^19] This corresponds to $u={a}^{2}\left(6\,c\,{\rm cn}\left(az+b,c\right){\rm dn}\left(az+b,c\right)+6\,{c}^{2}\,{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1\right)$ and $\hat{u}=a^{2}(6\,{c}^{2}{\rm sn}^{2}\left(az+b,c\right)-{c}^{2}-1)$. In particular, in the rational limit, $u=0$ or $u=12/z^{2}$, where $v=2/z$ or $v=-2/z$ respectively. $v=2ac\,{\rm sn}(az+b,c)$, where $a,b$ and $c$ are arbitrary constants. In general, to obtain an operator of the same type, several permutations and re-factorizations are required (see Appendix 1 for more details). Instead of studying these intermediate operations separately, we consider their result:
$\hat{L}=\hat{B}\hat{A},$ (79)
where $\hat{B}$ and $\hat{A}$ are of the form (77) with $v$ replaced by some $\hat{v}$:
$\hat{B}=\partial_{z}^{2}+\hat{v}\partial_{z}-\hat{v}^{\prime}-\frac{\hat{v}^{\prime\prime}}{\hat{v}},\quad\hat{A}=\partial_{z}-\hat{v}.$ (80)
We recall that in the Schrodinger operator case (cf. (30)), the Darboux transformation was the involution $A=\partial_{z}-v\leftrightarrow A^{*}=-(\partial_{z}+v)$, which corresponds to permutation of factors in the Schrodinger operator. This is, in fact, the involution $v\to-v$. Let us apply similar involution:
$\hat{v}=-v$
to our third-order operators. Then, from (77, 78), it follows that for such a transformation
$\hat{u}=u-6(\log\varkappa)^{\prime\prime}.$ (81)
Since $\hat{A}=\partial+v=\partial+\varkappa^{\prime}/\varkappa$, the function $\varkappa^{-1}$ is an eigenfunction of $\hat{L}=\hat{B}\hat{A}$ corresponding to the zero eigenvalue:
$\hat{L}\varkappa^{-1}=0.$
The complete zero-level transform $\hat{\varkappa}$ of $\varkappa$ is the general solution of equation
$\hat{L}\hat{\varkappa}=0.$ (82)
It can be obtained from its particular solution $\varkappa^{-1}$ by elementary methods: Fist, we rewrite (82) as
$\hat{\phi}^{\prime\prime}-\hat{u}\hat{\phi}=0,\quad\hat{\phi}=\hat{\varkappa}^{\prime}.$
One of the linearly independent solution of the above second order equation equals $(\varkappa^{-1})^{\prime}$. Then, its general solution is
$\hat{\phi}=C(\varkappa^{-1})^{\prime}\int\frac{dz}{\left((\varkappa^{-1})^{\prime}\right)^{2}},$ (83)
[…]
where the primitives include arbitrary constants of integration. Thus, we have found the zero-level Darboux transformations (81, 84) for the third-order operators of the type $\partial_{z}^{3}-u\partial_{z}$.
In the case, when $\varkappa$ corresponds to an equilibrium configuration and its transform $\hat{\varkappa}$ is factorizable, $\hat{\varkappa}$ also corresponds to an equilibrium configuration. This can be shown by arguments that are slight modifications of those of section 5. Specifically, we need to replace $\psi$, $\hat{\psi}$ from section 5 with [^20] Obviously, derivative of factorizable function is also factorizable. $\phi=\varkappa^{\prime}=\prod_{i}(z-z_{i})^{{\cal Q}_{i}}$, $\hat{\phi}=\hat{\varkappa}^{\prime}=\prod_{i}(z-\hat{z}_{i})^{\hat{\cal Q}_{i}}$, and replace the Darboux transformation for potential (52) with the transformation (81).
[…]
$\hat{\varkappa}=q_{+}/q,\quad\varkappa=q/q_{-},$ (85)
where $q_{+}$, $q$ and $q_{-}$ are some functions of $z$, we write
$\hat{\varkappa}^{\prime}=p/q^{2},\quad\varkappa^{\prime}=p_{-}/q_{-}^{2},$ (86)
where
$p=q_{+}^{\prime}q-q_{+}q^{\prime},\quad p_{-}=q^{\prime}q_{-}-qq_{-}^{\prime}.$ (87)
Substituting (85,86) into transformation (84), or equivalently, substituting $\hat{\phi}=p/q^{2}$, $(\varkappa^{-1})^{\prime}=-\varkappa^{-2}p_{-}/q_{-}^{2}$ and $\varkappa=q/q_{-}$ into (83) and rewriting result in the differential form, we obtain
$q^{4}\propto p^{\prime}p_{-}-pp_{-}^{\prime}.$ (88)
Equations (87, 88) are, up to normalization factors, the first-order recurrence relations (25) with $q_{+}=q_{n+1}$, $q=q_{n}$, $q_{-}=q_{n-1}$ and $p=p_{n}$, $p_{-}=p_{n-1}$.
Therefore, the iterations $\varkappa_{n+1}=\hat{\varkappa}_{n}$ of the Darboux transformations (84) generate main, as well as terminating sequences in the $\Lambda=2$ case, where equilibrium positions of charges are zeros and singular points of $\varkappa_{n}^{\prime}$.
Closing this section, we note that the operators $L_{n}=(\partial_{z}^{2}-u_{n})\partial_{z}=\partial_{z}^{3}+6(\log q_{n})^{\prime\prime}\partial_{z}$ (see (72,74)), which correspond to the main $\Lambda=2$ sequences, are rational Lax operators of the Sawada-Kotera hierarchy. The functions $q_{n}$ are polynomial $\tau$-functions of the hierarchy. Such $\tau$-functions can be written in a Pfaffian form [27]. Details are provided in Appendix 1, where we also present intertwining operators for $L_{n}$. Looking ahead, we note that, in contrast to $\Lambda=1$ case, these intertwining operators cannot be used to obtain translating configurations. This is because they intertwine third-order operators, rather than second-order Schrodinger operators. We will discuss the implications of this fact in the conclusion section.
In the next section we will focus on translating configurations in the $\Lambda=1$ case.
@@ 10 Translating Configurations and Baker-Akhieser Functions
Translating configurations of vortices are solutions of (5). To generate them, we consider the following eigenfunction
$\psi(k,z)=C(k)\prod_{i=1}^{N}(z-z_{i}(k))^{{\cal Q}_{i}}e^{kz},$ (89)
of the Schrodinger operator $H$:
@@
The condition that the potential $u=\psi^{\prime\prime}/\psi-k^{2}$ does not have simple poles (i.e., it is of form (50)) is equivalent to the equilibrium conditions (5) for a translating configuration. This can be shown by arguments similar to those employed for the static configurations in Section 5.
In the case of the factorization chain (36) starting with the free Schrodinger operator, there exists an eigenfunction, called the “rational Baker-Akhieser function” [^23] In the theory of integrable hierarchies of non-linear PDEs, the Baker-Akhieser function is a solution of an infinite system of linear PDE’s. Associated integrable hierarchy arises as a condition of compatibility of the system. Here we deal with the Baker-Akhieser function of the KdV hierarchy., which has the form (89). It is an eigenfunction of $H_{j}$, which is obtained by the action of an intertwining operator with rational coefficients on the eigenfunction $e^{kz}$ of the free Schrodinger operator, i.e., $\Psi=C(k)T_{j}[e^{kz}]$. Since the $z$-independent factor $C(k)$ does not play role in our considerations, we will use the eigenfunction
$\psi_{j}=T_{j}[e^{kz}]=\frac{p_{j}}{q_{j}}e^{kz},$ (90)
where $p_{j}$ and $q_{j}$ are polynomials in $z$. As we have shown in section 6, the corresponding potential $u_{j}$, obtained by $j$ Darboux transformations from $u_{0}=0$, is of the form (50). Since $u_{j}$ does not contain simple poles, eigenfunction (90) corresponds to a translating configuration. It will be shown in section 14 that the class of such configurations includes all possible translating patterns consisting of finite number of vortices with ${\cal Q}_{i}=\pm 1$.
A constructive way to write down $\psi_{j}$ is to use the Wronskian representation of $T_{j}$, given by (45, 46). Then, from (90), we get
@@
$p_{j}=e^{-kz}{\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j},e^{kz}\right],\quad q_{j}={\cal W}\left[{\cal X}_{1},{\cal X}_{2},\dots,{\cal X}_{j}\right]=P_{j},$ (91)
where $P_{j}$ is the $j$th Adler-Moser polynomial. When $k\not=0$, the degree of $q_{j}(z)$ equals to that of $p_{j}(z)$. This corresponds to the neutrality condition (8) for an equilibrium in the homogeneous electric field $k\not=0$. In the $k\to 0$ limit
@@ 11 Equilibria in Non-Homogenous Background Fields
$\psi(z)=\prod_{i=1}^{N}(z-z_{i})^{{\cal Q}_{i}}e^{\Phi(z)},\quad H\psi=\lambda\psi,$ (92)
@@
Equations (93) are the equilibrium conditions for charges in an external harmonic electric field [^24] Such conditions were first presented by Stieltjes [45, 46] for roots of orthogonal polynomials., or equilibrium of vortices in an irrotational background flow, see Section 1. The function $\Phi(z)$ is a potential of the external field/background flow. Provided (93) holds
@@
$\tilde{\psi}(z)=\prod_{i=1}^{\tilde{N}}(z-\tilde{z}_{i})^{{\cal\tilde{Q}}_{i}}e^{\Phi(z)},\quad H\tilde{\psi}=\eta\tilde{\psi},$
@@
will be also quasi-rational of the form similar to (92), i.e.,
@@
does not have first-order poles, $\hat{\psi}$ corresponds to a new equilibrium configuration in the same external field.
Thus, we can produce various equilibrium configurations by applying sequence of Darboux transformations to a set of eigenfunctions of $H$ of the form
$\varkappa_{i}=R_{i}(z)e^{\Phi(z)},\quad H\varkappa_{i}=\lambda_{i}\varkappa_{i},$
where $R_{i}(z)$ are rational in $z$. According to the Darboux-Crum theorem (39,40), the results of compositions of $n$ transformations can be written in the Wronskian form
$\psi^{(n)}(z)=\frac{{\cal W}[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n}}]}{{\cal W}[\varkappa_{k_{1}},\varkappa_{k_{2}},\dots,\varkappa_{k_{n-1}}]}.$ (95)
For example, for a harmonic oscillator $u=z^{2}/2$, $\Phi(z)=-z^{2}/2$, and $R_{n}(z)$ is the $(n-1)$th Hermite polynomial. In this case, transforms (95) correspond to various equilibria of vortices/charges in the quadrupole background flow/external field (for details and other examples, see e.g. [19, 32]). Note that, in contrast to the systems considered in the previous sections, configurations corresponding to Hermite polynomials do not depend non-trivially on continuous parameters and are rather parametrised by sequences of distinct integers $0<k_{1}<k_{2}<\dots<k_{n-1}$. These correspond to $\eta_{0}=\lambda_{k_{1}},\eta_{1}=\lambda_{k_{2}},\dots,\eta_{n-1}=\lambda_{k_{n}}$ in the chain (35).
@@ 12 Equilibria of Vortex Streets - Odd Family
A periodic vortex street is a vortex pattern consisting of an infinite number of vortices situated periodically along some direction. A vortex street can be equivalently considered as a configuration of $N$ vortices (or charges) on a (flat) cylinder. In what follows, the terms “vortex street”, “periodic configuration” or “periodic pattern” refer to a configuration of a finite number of charges on cylinder. Without loss of generality, we can set the translation period to $\pi$. Since $\sum_{n=-\infty}^{\infty}\frac{1}{z+\pi n}=\cot z$, the stationarity condition (5) for the translating (with a speed ${\rm i}\bar{k}/(2\pi)$) vortex street becomes
@@
and for static periodic patterns (see (6)), we have
@@
Similarly to vortex patterns on the plane, which are related to the rational solutions of the KdV hierarchy or the odd bispectral family, the trigonometric soliton solutions of the hierarchy are related to the patterns on a cylinder [24, 28, 32]. This can be demonstrated by methods similar to those employed in Sections 5, 10.
@@
where ${\cal Q}_{i}$, $z_{i}$ correspond to the periodic static configuration (97). The “trigonometric-rational” here means that all ${\cal Q}_{i}$ are integers. Trigonometric-rational functions are special cases of the trigonometric-factorizable functions, where ${\cal Q}_{i}$ are real. Function (98) is an eigenfunction of the Schrodinger operator
@@
similarly to the rational case, we come to the conclusion that the static equilibrium condition is equivalent to the absence of first-order poles in the potential and
@@
A family of vortex street configurations can be generated by the chain of Darboux transformations (35) with $\eta_{i-1}=k_{i}^{2}$, where $0<k_{1}<k_{2}<\dots<k_{i}<\dots$ is a sequence of positive integers. The chain begins with the free Schrodinger operator $H_{0}=-\partial_{z}^{2}$, and is constructed using “seed functions” that are its periodic eigenfunctions:
@@
Then, according to the Crum theorem (39,40), at $n$th step of the chain we have
@@
Clearly, $\tau_{n}(z)$ are trigonometric polynomials, $\psi^{(n)}(z)$ are trigonometric-rational functions and corresponding potentials
@@
are of the form (99) and do not have simple poles. Thus, the eigenfunctions (101) correspond to static vortex street configurations. Note that the potentials (102) are periodic $n$-soliton solutions of the KdV hierarchy, and $\tau_{n}$ are its $n$-soliton $\tau$-functions (see, e.g., [37]). We will demonstrate in Section 14 that this set of configurations actually includes all static vortex street patterns with ${\cal Q}_{i}=\pm 1$.
For translating configurations (see (96)) related to the soliton solutions, we can apply method similar to that used in section 10. By analogy with quasi-rational (89), we now consider the trigonometric quasi-rational eigenfunctions:
@@
According to the Darboux-Crum theorem (39), the eigenfunctions obtained through a sequence of the Darboux transformation from the eigenfunction $e^{kz}$ of the free Schrodinger operator are given by:
@@
where $\varkappa_{k_{i}}$ are given by (100). Obviously, $\psi^{(n)}(k,z)$ are of the form (103) and, therefore, correspond to the translating vortex street configurations. Since each $\varkappa_{k_{i}}$ (100) depends on a free parameter, both the translating and static configurations related to soliton solutions depend on $n$ free parameters as well as on the sequence of integers $0<k_{1}<k_{2}\dots<k_{n}$. Up to a $z$-independent common factor $C(k)$, (104) is the trigonometric Baker-Akhieser function of the KdV hierarchy.
The class of configurations determined by (104,100) includes all possible translating vortex street patterns with ${\cal Q}_{i}=\pm 1$, as will be shown in Section 14.